### v1

# Neutral Barium in Solid Neon: Optical Spectroscopy and First Excited State Lifetime

Abstract
Matrix isolation spectroscopy enables probing atomic properties in controlled cryogenic environments. Here we present a spectroscopic study on neutral barium atoms embedded in a neon cryogenic crystal at 6.8 K, extending previous investigations performed in other noble gas hosts.
The visible and near-infrared emission spectra were recorded under two different laser excitation schemes. First, 10-ns laser pulses at 355 nm were used to directly excite high-lying energy levels of barium, enabling the observation of fluorescence cascades. Second, a tunable continuous-wave laser operating between 700 nm and 900 nm allowed us to determine the matrix-induced shifts of barium energy levels relative to their vacuum values, as well as the inhomogeneous linewidths of the observed transitions and to perform lifetime measurements.
Our results confirm multiple radiative pathways and matrix-induced relaxation channels affecting the 5d6s and 6s6p barium manifolds. Furthermore, we present the first lifetime measurement of the barium 5d6s;3D1 state in a neon crystal, yielding 0.39 $\pm$ 0.02 s, with a predicted increase of about 10% at 2 K.
The study of fluorescence and spectroscopic properties of barium isolated in neon represents an important step toward future searches for the electron electric dipole moment using barium monofluoride in neon matrices, where neutral barium atoms may act as unavoidable impurities and potential sources of background and systematic limitations.
preprint: AIP/123-QED
I Introduction
In recent years, spectroscopy of atoms and molecules embedded into noble gas cryogenic crystals has developed significantly, as it plays a key role for high precision measurements within the matrix isolation technique (MIT). MIT enables the confinement of highly reactive species (dopants) within an inert, low‑temperature host DUBEY2018317, to investigate their characteristics and behavior while interrogating them at a fixed point in space and simultaneously achieving a high density of dopants in the crystal Weinstein2005.
Other techniques to investigate atoms or molecules include slowing a production beam Patterson2007 or trapping them through Magneto Optical Traps (MOTs) Raab1987. These methods enable in-gas investigation of the species, at the cost of the number of molecules. However, for fundamental physics research, having a large number of atoms or molecules enhances the sensitivity of the measurements Safronova2018. Therefore, in this study, we discuss the production of an inert cryogenic environment with a high density of dopants and investigate them through spectroscopic measurements.
The investigation method used is Laser Induced Fluorescence (LIF) as it allows for high signal to noise ratio and good sensitivity for atomic or molecular spectroscopy Kinsey1977.
Neutral barium (Ba) has a favorable electronic structure to be observed in noble gas matrices Callear1973, Trenary1973, and has been recently characterized in matrices such as Ar, Kr and Xe using LIF spectroscopy, revealing multiple occupancy sites and Jahn-Teller splitting McCaffrey, McCaffrey_BaSr_inRGmatrices. Further studies in solid helium have also provide new insights on how barium, under very low perturbation conditions, shows a very long lifetime of atomic metastable states of the orders of seconds Weis_BaHelium.
Figure 1: Top view sketch of the experimental setup including production of a gas phase barium beam, crystal growth, excitation, and fluorescence detection: A) Barium production and doped neon crystal growth setup: two lasers are simultaneously used. An Nd:YAG laser (light blue arrows in the figure) is used for the ablation of the metallic barium target. An Avesta laser (red arrows in the figure) at 12636 cm-1 (791.13 nm) is used to probe the presence of barium atoms, ensuring efficient production both inside and outside the cell. B) Excitation and fluorescence detection scheme: the gas composed of Ba and neon is sprayed onto the CaF2 substrate (orange rectangle) at 6.8 K, which is permanently tilted to minimize the collection of excitation light in the fluorescence detection system. Purple arrows indicate the excitation laser. A custom built lens assembly allows us to fix the first collection lens at about 30 mm from the crystal growth spot, enabling efficient collection of the emitted signal which is then focused on either a spectrometer or a photodiode, depending on the measurement.
In this paper, we extend the approach conducted in previously cited works, proving the first successful growth of good quality Ba doped neon crystals at temperatures ranging from 6.8 to 9 K. Among the hosts previously studied, neon is a promising matrix for MIT, as it can offer good optical transparency and reduced perturbation of embedded species, as proven in previous studies of atoms embedded in cryogenic neon crystals 1Yb_Neon_Lambo, dargyte2021optical_Rb_Neon_Spin, braggio2022spectroscopy_RbNEon. Therefore, this approach enables direct comparison of Ba properties in our neon crystal lattice with respect to the previously studied ones.
In this work, we investigate neutral barium atoms embedded in a neon matrix, where neon serves both as a buffer gas in our atomic source and as a matrix deposition gas for crystal growth. This systems enables LIF spectroscopic measurements of barium atoms in the crystal, using single and double laser excitation in either continuous or pulsed mode. These methods enable selective excitation of barium energy levels. We also measure, for the first time, the lifetime of the metastable state 5d6s 3D1 of barium in a neon matrix.
This approach creates not only a new base for spectroscopic analysis of barium in inert cryogenic matrices, but also an essential experimental framework for embedding molecules, such as barium monofluoride (BaF), in cryogenic crystals. BaF is a molecule with great characteristics for low energy, high precision measurements such as the electron electric dipole moment (eEDM) through the MIT.
BaF production will include deposition of some neutral barium presence in the matrix, unless it is filtered out yau2024specularreflectionpolarmolecules, corriveau2024matrixisolatedbariummonofluoride at the cost of a loss in the number of molecules. Therefore, understanding the spectral signatures and dynamics of Ba will be essential, as it will provide a crucial spectroscopic baseline for future BaF matrix isolation measurements in a cryogenic inert environment such as neon Li_2023, corriveau2024matrixisolatedbariummonofluoride or parahydrogen Messineo_pH, Borghesani_2023, as recently proposed by our group within the DOCET experiment.
II Experimental setup
The experimental procedure of our spectroscopic work consists of three main stages: the growth of a neon cryogenic crystal with embedded Ba atoms and LIF detection outside the vacuum chamber.
II.1 Barium doped neon crystals production
[…]
Ablation laser pulses at 1064 nm with a repetition rate of 10 Hz and energies between 10 and 30 mJ, generated from a EKSPLA NL300 Nd:YAG laser, hit the metallic barium target, fixed on one side of the cell, on a focused spot with diameter of approximately 0.5 mm. The plasma plume evaporates Ba atoms, carried out of the cell by neon flowing at 5–20 SCCM through a 3 mm aperture. The ablation spot is periodically shifted to maintain stable production.
To probe the production of neutral barium atoms in the gas phase, before deposition, two absorption measurements are performed simultaneously, as already done in previous studies for BaF mooij2024novel, inside and outside of the cell. An Avesta Ti:Sa laser is used, with a linewidth of approximately 2 GHz, tunable over a 14 300 - 11100 cm-1 range (700 - 900 nm). The laser is tuned to 12 636 cm-1 (791.13 nm), probing the transition from the ground state 6s2 $^{1}\text{S}_{0}$ to 6s6p $3\text{P}_{1}$. The barium energy levels used in this work are shown in Figure 2. An absorption measurement probes the production directly inside the cell, where the laser is sent through a window at the back of the cell and exits from the extraction hole towards a Thorlabs DET36A/M silicon photodiode. The second method consist in a differential measurement with a Thorlabs PDB210A/M. To do so, a laser beam is equally split, one part is used as reference signal, while the other passes 5 mm from the extraction hole, in a dual pass configuration, aligned through two diaphragms at the side of the extraction hole. Noise is highly suppressed in this configuration, allowing for the detection of bunches of atoms coming out of the source every ablation pulse.
Crystal growth happens on a CaF2 substrate mounted on a copper holder that is connected to a two-stage helium pulse tube cryocooler Sumitomo RP62B, maintaining a temperature of 6.8 K. A heater is positioned on top of the cold finger, allowing gradual temperature ramps and, around 16 K, full sublimation of the neon crystal and its dopants.
Following the production of Ba atoms, the neon buffer gas carries them to the CaF2 substrate, where the crystal grows for 20 to 60 minutes. Before initiating the ablation process, a 10 SCCM neon gas flow is maintained for 10 minutes, allowing optimal base growth of a pure neon crystal and ensuring better sublimation of the doped crystal.
II.2 Optical setup for fluorescence collection
Fluorescence measurements are performed using multiple laser systems, hitting the doped crystal positioned at 45 degrees with respect to the incoming barium beam, as shown in Figure 1 B). A combination of two identical Avesta Ti:Sa laser is used to deliver 10–200 mW on a 1 mm wide spot size on the crystal surface for single and double laser excitation, as described in Section III.2. Pulses of 28 195 cm-1 (354 nm) are generated through the third harmonic of a 1064 nm Nd:YAG laser. At a repetition of 10 Hz, the pulses are 10 ns long and on a range between 1 to 2.5 mJ and are used to excite neutral barium to high energy states, as describe in the next section.
Light collection occurs almost orthogonally from the excitation beam, to reduce the unwanted excitation signal and optimize the fluorescence detection.
The light detection system consisted of a two lens assembly to collect fluorescence emitted from the substrate. The lenses are mounted inside a custom-built cylindrical holder that extended into the cryogenic chamber, allowing us to set the first lens at a distance of approximately 30 mm from the crystal surface, as shown in Figure 1. To isolate the fluorescence signal from the excitation laser light, spectral filters are used, depending on the measurement done. Detection of the fluorescence light occurred through an Hamamatsu InGaAs photodiode for lifetime measurements, or two spectrometers: Spectral Products SM642 for the visible range and Ocean Optics NIR Quest InGaAs for near-infrared (NIR) measurements of the first excited state of barium.
III Results and Discussion
In this section, we present the spectroscopic characterization of neutral barium atoms isolated in a neon matrix at 6.8 K. We first analyze the high-energy excitation regime using pulsed third harmonic (355 nm) laser light, followed by the investigation of low lying transitions probed through narrow linewidth continuous wave lasers in the 700 - 900 nm range. Finally, we report the lifetime measurement of the metastable 5d6s 3D1 state in this environment.
III.1 High energy excitation of barium in neon matrix
[…]
The work of Lebedev et al. Weis_BaHelium demonstrate that a third harmonic of an Nd:YAG laser, with pulses at 28 195 cm-1 (355 nm) can excite a wide range of barium states by both radiative or non-radiative transitions in a solid helium matrix. Among these, long lived metastable state of barium, such as 5d6s 1D2 and the 5d6s 3Dj ($j=1,2,3$) manifold, act as an intermediate reservoir, and when a second pulse excites them, they can reach excited state at energies over 30 000 cm-1 with respect to the energy of the ground state. This method leads to barium fluorescence emission through a radiative cascade, as from the highest excited state the atom can populate a metastable state before decaying into the ground state.
We adapted this approach to detect barium fluorescence in a neon crystal (vis-NIR), enabling the first comparison up to date with the work of Lebedev et al. Weis_BaHelium in solid helium.
As illustrated in Figure 3 and summarized in Table 1 we observe a broad set of barium fluorescence peaks when a third harmonic of an Nd:YAG pulsed laser is shined onto the crystal. The reported spectra is obtained by subtracting the fluorescence spectra of the CaF2 substrate from the one obtained from the barium doped neon matrix on the substrate.
Figure 3: Fluorescence spectrum of atomic barium in a neon cryogenic matrix at 6.8 K when excited with a third harmonic (355 nm) of a Nd:YAG laser with a repetition rate of 10 Hz. In gray: experimental data (baseline-subtracted); black solid: multi-Gaussian total fit; black dashed: individual fitted transitions. The region above 750 nm is magnified ×5. List of cascade fluorescence detected: A - 6s9s $^{3}\text{S}_{1}$ to 6s6p $^{1}\text{P}_{1}$; B - 6s6d $^{3}\text{D}_{1}$ to 6s6p $^{3}\text{P}_{0}$; C - 6s6p $^{1}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$; D - 6s6d $^{3}\text{D}_{3}$ to 6s6p $^{3}\text{P}_{2}$; E - 5d6p $^{3}\text{F}_{2}$ to 5d6s $^{3}\text{D}_{1}$; F - 5d6p $^{3}\text{F}_{2}$ to 5d6s $^{3}\text{D}_{2}$; G - 6s6p $^{3}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$; H - 5d6s $^{1}\text{D}_{2}$ to 6s2 $^{1}\text{S}_{0}$.
Table 1: Barium emission transitions in a neon crystal at 6.8 K were observed following excitation with the third harmonic (355 nm) of a Nd:YAG laser. The spectral shifts are referenced against the corresponding free atomic barium lines, listed in the lower portion of the table. Transition H is an electric dipole forbidden line in the gas phase NIST_ASD. Symbol $\ast$ = multiple transitions involved.
Observed LIF peaks
Label $\lambda$ (nm) $\nu$} (cm-1) Shift $\delta$ (cm-1) FWHM $\Delta\nu$ (cm-1)
A 530.3 18 854 +12+12 330
B 539.0 18 551 +122+122 354
C 556.5 17 968 93-93 156
D 573.8 17 427 +124+124 411
E 768.5 13 011 18-18 88
F 778.9 12 838 9-9 113
G 793.7 12 599 35-35 117
H 875.5 11 422 +27+27 102
Corresponding free atom transitions
Label $\lambda$ (nm) $\nu$} (cm-1) Free gas transition
A 530.6 18 842 6s9s 3S1 - 6s6p $^{1}\text{P}_{1}$
B 542.5 18 429 6s6d 3D1 - 6s6p $^{3}\text{P}_{0}$
[…]
D $\ast$ 577.9 17 303 6s6d $^{3}\text{D}_{3} -$ 6s6p $^{3}\text{P}_{2}$
E 767.2 13 030 5d6p $^{3}\text{F}_{2} -$ 5d6s $^{3}\text{D}_{1}$
F 778.0 12 848 5d6p $^{3}\text{F}_{2} -$ 5d6s $^{3}\text{D}_{2}$
[…]
We identified a total of eight fluorescence peaks. Comparing our data with the spectroscopic results of Lebedev et al. on barium in solid helium Weis_BaHelium, four fluorescence lines match those previously reported.
In general, the relatively narrow wavelength shifts with respect to the free barium atomic lines and the relatively small full width at half maximum (FWHM) allow us to reliably identify five electronic state transitions (C, E, F, G, H, see Figure 3), as shown in Table 1. Among these, the largest observed shift is for peak C (93-93 cm-1), while the smallest is for peak F (9-9 cm-1). The magnitude of these shifts is smaller than those reported for barium isolated in rare gas matrices McCaffrey.
However, some emission features may not be fully resolved because of their relatively large FWHM. Peaks A and B are identified through a multi-Gaussian fit of the same broad feature, as shown in Figure 3, and are identified with small shifts of +12+12 and +122+122 cm-1, respectively, with respect to the one measured in free atomic Ba. Peak D, centered at 17 427 cm-1 (573.8 nm), might arise from a convolution of multiple energy levels. In this wavelength range, transitions such as 6s6d $^{3}\text{D}_{1,2,3}$ to 6s6p $^{3}\text{P}_{1}$ (17 091, 17 236, and 17 303 cm-1), or 6p2 $^{3}\text{P}_{2}$ to 6s6p $^{1}\text{P}_{1}$ (17 284 cm-1), could all contribute to the observed line, explaining its large FWHM (411 cm-1). The reported shift in Table 1 is referenced to the 17 303 cm-1 transition.
The FWHM values can be directly compared with those obtained by Lebedev et al. Weis_BaHelium of barium in solid helium at 1.5 K. Peak C, corresponding to the 6s6p $^{1}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$ transition, has a width of $\Delta\nu=156\ \text{cm}^{-1}$, broader than the helium value ($\delta\nu=30\ \text{cm}^{-1}$). However, peak E, F and G show the same width with respect to the same transitions observed in solid helium.
Overall, both the spectral shifts and FWHM measured here are consistent with those reported in solid helium.
III.2 Excitation with single and double laser
In this section, our objective is to investigate the doped crystal using continuous-wave lasers, following the work of Li et al. Li_2023. To this end, we irradiated 1 mm-wide spots on the crystal surface with Ti:Sa lasers delivering up to 100 mW, and measured excitation spectra by tuning the laser over its operational range of 14 300–11 100 cm-1 (700–870 nm).
Single laser excitation
Figure 4: Energy level diagram (not to scale) of the single laser pumping scheme of neutral barium atoms in a neon matrix. Both excitation lasers, at 13 514 cm-1 (749.8 nm, green) and 13 369 cm-1 (740.5 nm, red), populate the 6s6p 3Pj manifold. The system is relaxed to the states 5d6s 3D1 and 5d6s 1D2, which then decay to the ground state. These population transitions are detected respectively at 9 033.4 cm-1 (1107.1 nm, green) and 11 494.2 cm-1 (870.0 nm, red).
[…]
These excitation wavelengths are far detuned from direct transitions and likely populate the 6s6p 3Pj manifold through matrix-induced relaxation, as shown in Figure 4. Those states can decay to the ground state through radiative cascade involving the 5d6s 3Di (i = 1,2,3) and 1D2 states. Direct fluorescence from the 3Pj states is not detected, as it spectrally overlaps with the excitation light. Our working hypothesis is that the excitation laser directly addresses the 6s6p 3P2 level at 13 514.7 cm-1 (740.0 nm), a strongly suppressed transition in the free atom. However, in the matrix environment, the local perturbation breaks the atomic symmetries that govern the selection rules in the gas phase moroshkin2008atomic. As a result, transitions that are forbidden in free barium become weakly allowed, allowing the population transfer to the 6s6p 3P2 state and the subsequent relaxation into the 5d6s 3Di manifold or 5d6s 1D2 state.
As stated in precedent spectroscopy studies of Ba atoms in RG gas matrices, the line broadening of excitation is larger with respect to the fluorescence emission. In this case, the effect is likely enhanced by the indirect excitation discussed above.
Figure 5: 2D excitation–emission spectrum of barium in a neon crystal at 6.8 K when shined with a single laser at 35 mW on a 1 mm wide spot. A strong emission band is observed, with a maximum intensity at 11 415 cm-1 (876.0 nm), corresponding to population of the 5d6s 1D2 state (+20+20 cm-1 shift), when the laser is tuned at 13 504.3 cm-1 (740.5 nm). A long blue tail is visible until the reach of our laser at 700 nm of excitation. The residual signal of the excitation laser appears at longer wavelengths due to the reduced efficiency of the optical filtering.
Figure 6: 2D excitation–emission spectrum of barium in a neon crystal at 6.8 K when shined with a single laser at 35 mW on a 1 mm wide spot. An emission band is observed around 9033.4 cm-1 (1107.1 nm), corresponding to population of the 5d6s 3D1 state, with maximum intensity under excitation at 13 514 cm-1 (749.8 nm) and long blue tail of approximately 40 nm, similar to the one observed previously in Figure 5.
Interestingly, we also detected green fluorescence around 18000 cm-1 corresponding to the 6s6p 1P1 state, as shown in Figure 7 (a).
The photon energies provided by our excitation laser are insufficient to directly excite the 6s6p 1P1 state. Our hypothesis, further investigated in the following sections, is that since the photon energy of the Ti:Sa laser is insufficient to reach the 6s6p,1P1 state directly, the observed green fluorescence likely arises from multi step population of higher manifolds even under single frequency continuous excitation. This mechanism is analogous to the excitation scheme achieved with the 3rd harmonic pulsed laser, as discussed in Section III.1.
Figure 7: 2D excitation–emission spectra of barium atoms embedded in a neon crystal at 6.8 K, showing fluorescence in the green region. (a) Excitation with a single narrow linewidth laser at low energy allows population transfer from the first excited manifold to higher-lying states, whose fluorescence cascade produces green light comparable to that observed under third-harmonic excitation. (b) When a second identical laser is spatially overlapped with the first and tuned to 748 nm, which populates the 5d6s 3D1 state, additional population transfer is observed. New fluorescence pathways emerge, revealing alternative excitation channels to higher-lying states
Double laser excitation scheme and results
To test the feasibility of pumping schemes with continuous lasers, we performed further measurements using a combination of two identical Ti:Sa laser.
Since our strongest source of fluorescence signal is the green one, we first focused on it. The first laser is fixed at the peak of population transfer in the 5d6s 3D1 state (9 033.4 cm-1), while we scanned with the second excitation laser over the same frequency region as before (14 300 - 11 100 cm-1), overlapping the two laser interaction regions in the crystal at a power of 40 mW.
The resulting spectrum, shown in Figure 7 (b), reveals an increase in the fluorescence intensity of the previously detected signal, as well as the appearance of new excitation combinations, resulting in high intensity fluorescence of the 6s6p 1P1 state.
Interestingly the 6s6p 3P1 state exhibits two dominant fluorescence bands centered at 18 622 (557 nm) and 17 967 (538 nm) cm-1. Similar features have been extensively studied in Ar, Kr, and Xe matrices McCaffrey, where they are attributed to the influence of lattice site occupancy on barium emission properties.
Figure 8: Energy level diagram (not to scale) of the proposed double laser pumping scheme for barium in a neon matrix. The first laser (red arrow) at 13 369 cm-1 (749 nm) indirectly populates the 5d6s 3D2 state. From this level, a second laser (blue arrow) tuned at 12 849 cm-1 (780 nm) excites the atom to the higher 5d6p 3F2 state. According to the cascade dynamics reported in previous works Weis_BaHelium, this state decays to the 5d6s 1D2 level with emission at 10 670 cm-1 (937 nm), followed by a second fluorescence decay to the ground state 6s2 1S0 at 11 394 cm-1 (877 nm). The black arrows represent spontaneous emission pathways, while red arrows denote laser driven excitations. Dashed lines indicate indirect population routes from the 6s6p 3Pj manifold.
We observe a high intensity fluorescence emission when the excitation laser is tuned to approximately 14 180 cm-1 (705 nm). However, it does not correspond to any neutral atomic barium state, and no direct excitation pathway from the ground or low lying metastable states to higher lying electronic states has been identified at this wavelength. This discrepancy suggests that the observed fluorescence may arise from matrix induced perturbations, site selective effects, or indirect population mechanisms through long lived intermediate states. Further investigation is beyond the scope of this work.
To further test the feasibility of the double–laser excitation, we designed a pumping scheme path that populates 5d6p 3F2 at 22 065 cm-1 (453 nm) only if we strongly populate the manifold 6s6d $^{3}\text{D}_{j}$, as shown in Figure 8. Although the excitation bands are $\approx$1000 cm-1 wide, our measurements demonstrate that this scheme works exclusively via indirect population of the first manifold, specifically through the 5d6s 3D2 state. A second laser, set at 12 849 cm-1, can then couple this state and populate the 5d6p 3F2 state laying at 22 064.622\,064.6 cm-1 (453 nm). This scheme relays on the evidence shown earlier that excitation near 13 369 cm-1 (748 nm) populates the 5d6s 3D1 and hypothetically the other states of the same 5d6s 3Dj manifold.
Figure 9: 2D excitation–emission spectra of the double-laser test probing the 5d6p;3F2 fluorescence cascade of barium in a neon matrix. (a) Two spatially overlapped continuous-wave beams (40 mW each) illuminate the crystal: the first, fixed at 13 369 cm-1 (749 nm), populates the first excited-state manifold (5d6s 3D2), while the second is scanned near 12 849 cm-1 (780 nm) to drive the 5d6s 3D2 to 5d6p 3F2 transition. A distinct fluorescence band at 10 670 cm-1 (937 nm) appears, together with an enhanced 11 395 cm-1 (877 nm) band, consistent with the cascade 5d6p 3F2 to 5d6s 1D2 (937 nm) followed by 5d6s 1D2 to the ground state (870 nm). (b) When the first laser is tuned to 740 nm, the 937 nm band disappears while the 877 nm baseline remains, confirming that the previous excitation scheme selectively populates the 5d6s;3D2 state, which is required to reach 5d6p 3F2.
The expected result of the pumping scheme is a cascade fluorescence, starting from the level 5d6p 3F2 and decaying to 1D2 at 10 670 cm-1 (937 nm), followed by an emission to the ground state at 11 395 cm-1 (878 nm), as indicated by the straight black lines in Figure 8.
Experimentally, the first laser is fixed at 13 369 cm-1, while the second is scanned between 13 300 and 12 700 cm-1. The result, shown in Figure 9 (a), reveals expected LIF emissions at 10 640 cm-1 (939 nm) along with an enhanced line at approximately 11395 cm-1 (878 nm), but only when the second laser is tuned around 12 805 cm-1 (781 nm), close to the expected wavelength, shifted by just 44 cm-1. This observed fluorescence confirms the predicted cascade behavior from 5d6p 3F2 to 5d6s 1D2 and then to the ground state.
To further test the scheme, if we tune the first laser to 13 504 cm-1 (740 nm), a known population maximum of the 5d6s 1D2 state, no excitation pathway to the 5d6p 3F2 level is observed, as shown in Figure 9 (b), probing our selective pumping scheme functioning.
III.3 Lifetime of first excited state 5d6s 3D1
As shown in the previous sections, and schematically in Figure 4, a single laser allows us to populate the 5d6s 3Dj manifold, similar to the earlier work of Lebedev et al. Weis_BaHelium, where the 5d6s 3D1 state of barium in a helium matrix at 1.5 K is populated using the pulsed second harmonic of a Nd:YAG laser at approximately 18 80018\,800 cm-1 .
In our experiments with barium in a solid neon at 6.8 K, this state is populated using continuous wave laser light tuned to 13 514 cm-1, enabling a direct comparison with the lifetime measurement performed in solid helium and in the vacuum gas phase.
The lifetime measurement is carried out using an Electro Optic Modulator (EOM), driven by a pulse generator. This configuration allowed us to modulate the excitation beam intensity, switching the laser on and off in 200 ns.
The decay of the state population from 5d6s 3D1 to the ground state is shown in Figure 10, as well as the fit to compute the measurement of the lifetime ($\tau$) following the model: $I=I_{0}e^{-t/\tau}+C$, where $I_{0}$ is the initial intensity and $C$ is the background intensity due to the non-zero extinction of the EOM. This background level is shown in Figure 10 as a gray dashed line.
The resulting lifetime of the 5d6s 3D1 state of barium in the neon matrix is found to be $0.39\pm 0.02$ s.
Figure 10: Fluorescence decay of the 5d6s 3D1 state of barium embedded in solid neon at 6.8 K. The red curve shows an exponential fit, yielding a lifetime determination of 0.39 $\pm$ 0.02 s. The gray dashed line represents the background level of fluorescence light due to imperfect extinction of the excitation light when the EOM is off. Inset: Fluorescence intensity dependence on the crystal temperature. At temperature higher than 8 K, intensity in the fluorescence shows a plateau.
We obtain a lifetime more than two orders of magnitude higher than the gas phase measurement, which is estimated to be 1.5 $\pm$ 0.2 ms AdamsLeiHusain+2001+1205_lifetimeBaGas. This increase might indicate the weak interaction of the host matrix, as neon low phonon density and weak interaction with the embedded atoms enable long lived excited states. For a comparison, in solid helium, an even better lifetime of $2.72\pm 0.04$ s is observed at a temperature of 1.5 K Weis_BaHelium.
Lifetime measurements have been conducted at different temperatures, as, ranging from 6.8 to 8 K, due to temperature limitations of the setup. A measurement of the LIF intensity depending on the substrate temperature is shown in the inset of Figure 10. At temperatures higher than 8 K a plateau is observed. This is an unknown phenomena, that might be due to the sublimation of the neon crystal starting at approximately 8 K in a vacuum of $10^{-7}$ mBar.
Figure 11: Temperature dependence of the lifetime of the 5d6s;3D1 state of barium embedded in solid neon. In red a fit of Eq. 1, the rate equation governing the radiative and non-radiative decay of an atom embedded in a cryogenic crystal lattice. From such a fit, a lifetime of 0.42 $\pm$ 0.03 s is expected at 2 K. The shaded red band represents the one $\sigma$ confidence interval of the fit.
A rate equation model has been used to fit the obtained lifetime and extrapolate the possible lifetime values at lower temperatures. It considers both radiative ($\tau_{r}$) and non-radiative ($\tau_{nr}$) decay channels, described by the two termsfit_description_PL, torVergata_fitArrhenius:
$\frac{1}{\tau(T)}\;=\;\frac{1}{\tau_{r}}\;+\;\frac{1}{\tau_{nr}}\,\exp\!\left(-\frac{E_{a}}{k_{B}T}\right)\,.$ (1)
Fit results are shown in Table 2, where $E_{a}$ is the activation energy . The lifetime is expected to increase to approximately 0.42 $\pm$ 0.03 s at 2 K.
Table 2: Fit parameters from the rate equation model describing the temperature dependent lifetime of the 5d6s 3D1 state of barium in solid neon.
Parameter Value
$\tau_{r}$ $0.42\pm 0.02$ s
$\tau_{nr}$ $0.16\pm 0.01$ s
$E_{a}$ $(1.4\pm 0.3)\times 10^{-3}$ eV
IV Conclusions
For the time in literature, we observed barium atoms embedded in a neon matrix through a combination of cascade florescence and selective excitation using a third harmonic 1064 nm pulsed laser and single or double continuous wave lasers, exploiting long living metastable state of barium within the matrix.
The observed fluorescence spectra display small shifts relative to the gas phase transitions and moderate inhomogeneous broadening, indicating the weak perturbation exerted by the neon matrix. The main 6s6p 1P1 to 6s2 1S0 transition was observed at 556.5 nm, with a shift with respect to the free gas phase of 93-93 cm-1 and a linewidth of 156 cm-1. Overall, all identified transitions have shown relatively small matrix shifts and widths with respect to heavier RG matrices (Ar, Kr, Xe). The 5d6p 3F2 to 5d6p 3Pj transitions show shifts comparable to those in helium, confirming neon’s low perturbation nature.
Using a single tunable laser between 700 and 900 nm, we identified indirect excitation pathways populating the 5d6s 3D1 and 1D2 metastable states via matrix induced symmetry breaking. With two simultaneously applied lasers, we demonstrated an excitation pathway from the 5d6s 3D2 state to 5d6p 3F2, followed by a cascade fluorescence sequence 5d6p 3F2 to 5d6s 1D2 to 6s2 1S0, confirming the feasibility of selective excitation within the matrix.
This spectroscopy work on neutral barium in solid neon provides a crucial benchmark for future BaF spectroscopy within the same cryogenic matrix. Since the two species share similar spectral regions, and barium is produced during the production of BaF, a detailed understanding of Ba optical transitions, line widths, and fluorescence dynamics is essential to correctly identify and separate atomic and molecular contributions in mixed Ba/BaF matrices.
[…]
Our lifetime measurement serves as a reference point for future studies aiming to explore similar systems in parahydrogen matrices. Previous work by Lancaster et al. Lancaster_2021_NevsPh found no detectable fluorescence for rubidium in parahydrogen, whereas clear emission was observed for boron Tam2000_boron_pH and other molecules Huang2016_molecules_pH. A similar investigation for Ba and BaF will be crucial to establish parahydrogen as a next-generation host for matrix-isolated spectroscopy and precision measurements of fundamental properties.
Acknowledgements.
We express our sincere gratitude to A. C. Vutha for contributing to insightful discussions that greatly contributed to this work.
We also thank E. Berto, F. Calaon, M. Tessaro, M. Rebeschini, and M. Zago for their invaluable technical support.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author contributions
Alessandro Lippi: Conceptualization (equal); Data curation (lead); Formal analysis (equal); Investigation (equal); Methodology (equal); Software (lead); Visualization (lead); Writing – original draft (lead); Writing – review & editing (equal).
Giovanni Carugno: Conceptualization (equal); Funding acquisition (lead); Methodology (equal); Project administration (lead); Resources (equal); Supervision (lead); Writing – review & editing (equal).
Roberto Calabrese: Funding acquisition (equal); Resources (equal).
Federico Chiossi: Validation (equal); Writing – review & editing (equal).
Marco Guarise: Formal analysis (equal); Funding acquisition (equal); Resources (equal); Supervision (equal); Validation (equal); Writing – review & editing (equal).
Madiha M. Makhdoom: Investigation (equal); Writing – review & editing (equal).
Giuseppe Messineo: Conceptualization (equal); Investigation (equal); Validation (equal); Writing – review & editing (equal).
Jacopo Pazzini: Funding acquisition (equal); Resources (equal); Writing – review & editing (equal).
Founding
This research was funded by the PNRR MUR project PE0000023 – NQSTI, financed by the European Union—Next Generation EU. This work was also supported in part by the Italian MUR “Departments of Excellence grant 2023–2027—Quantum Frontiers”, by the “Progetti di Ricerca di Rilevante Interesse Nazionale” (PRIN grant 20227F5W4N, PRIN 2022), and by INFN CSNV within the framework of DOCET experiment.
Data availability
The data supporting the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@
 In recent years, spectroscopy of atoms and molecules embedded into noble gas cryogenic crystals has developed significantly, as it plays a key role for high precision measurements within the matrix isolation technique (MIT).
-MIT enables the confinement of highly reactive species (dopants) within an inert, low‑temperature host DUBEY2018317, to investigate their characteristics and behavior while interrogating them at a fixed point in space and simultaneously achieving a high density of dopants in the crystal Weinstein2005.
-Other techniques to investigate atoms or molecules include slowing a production beam Patterson2007 or trapping them through Magneto Optical Traps (MOTs) Raab1987.
+MIT enables the confinement of highly reactive species (dopants) within an inert, low‑temperature host Dubey et al.
+[2018], to investigate their characteristics and behavior while interrogating them at a fixed point in space and simultaneously achieving a high density of dopants in the crystal Upadhyay et al.
+[2020].
+Other techniques to investigate atoms or molecules include slowing a production beam Patterson and Doyle [2007] or trapping them through Magneto Optical Traps (MOTs) Raab et al.
+[1987].
 These methods enable in-gas investigation of the species, at the cost of the number of molecules.
-However, for fundamental physics research, having a large number of atoms or molecules enhances the sensitivity of the measurements Safronova2018.
-Therefore, in this study, we discuss the production of an inert cryogenic environment with a high density of dopants and investigate them through spectroscopic measurements.
-The investigation method used is Laser Induced Fluorescence (LIF) as it allows for high signal to noise ratio and good sensitivity for atomic or molecular spectroscopy Kinsey1977.
-Neutral barium (Ba) has a favorable electronic structure to be observed in noble gas matrices Callear1973, Trenary1973, and has been recently characterized in matrices such as Ar, Kr and Xe using LIF spectroscopy, revealing multiple occupancy sites and Jahn-Teller splitting McCaffrey, McCaffrey_BaSr_inRGmatrices.
-Further studies in solid helium have also provide new insights on how barium, under very low perturbation conditions, shows a very long lifetime of atomic metastable states of the orders of seconds Weis_BaHelium.
-Figure 1: Top view sketch of the experimental setup including production of a gas phase barium beam, crystal growth, excitation, and fluorescence detection: A) Barium production and doped neon crystal growth setup: two lasers are simultaneously used.
-An Nd:YAG laser (light blue arrows in the figure) is used for the ablation of the metallic barium target.
-An Avesta laser (red arrows in the figure) at 12636 cm-1 (791.13 nm) is used to probe the presence of barium atoms, ensuring efficient production both inside and outside the cell.
-B) Excitation and fluorescence detection scheme: the gas composed of Ba and neon is sprayed onto the CaF2 substrate (orange rectangle) at 6.8 K, which is permanently tilted to minimize the collection of excitation light in the fluorescence detection system.
-Purple arrows indicate the excitation laser.
-A custom built lens assembly allows us to fix the first collection lens at about 30 mm from the crystal growth spot, enabling efficient collection of the emitted signal which is then focused on either a spectrometer or a photodiode, depending on the measurement.
+However, for fundamental physics research, having a large number of atoms or molecules enhances the sensitivity of the measurements Safronova et al.
+[2018].
+Therefore, in this study, we discuss the production of an inert cryogenic environment with a high density of barium (Ba) atoms as dopants to investigate them through spectroscopic measurements as a testbench for future studies of more complex molecules such as barium monofluoride (BaF).
+The investigation method used is Laser Induced Fluorescence (LIF) as it allows for high signal to noise ratio and good sensitivity for atomic or molecular spectroscopy Kinsey [1977].
+Neutral barium has a favorable electronic structure to be observed in noble gas matrices Callear and Metcalfe [1973], Trenary and Shriver [1973], and has been recently characterized in matrices such as Ar, Kr and Xe using LIF spectroscopy, revealing multiple occupancy sites and Jahn-Teller splitting Davis and McCaffrey [2018, 2016].
+Further studies in solid helium have also provide new insights on how barium, under very low perturbation conditions, shows a very long lifetime of atomic metastable states of the orders of seconds Lebedev et al.
+[2011].
+Figure 1: Top view sketch of the experimental setup including production of a gas phase barium beam, crystal growth, excitation, and fluorescence detection: A) Ba doped neon crystal growth setup: two lasers are simultaneously used.
+An Nd:YAG laser (light blue arrows in the figure) is used for the ablation of the metallic barium target, which gets extracted from the cell through a hole in front of a CaF2 substrate.
+An Avesta laser (red arrows in the figure) at 12636 cm-1 (791.13 nm) is used to ensure efficient production both inside and outside the cell during ablation.
+B) Excitation and fluorescence detection scheme: the gas composed of Ba and neon is sprayed onto the CaF2 substrate (orange rectangle) at 6.8 K, which is permanently tilted to minimize the collection of excitation light (purple arrows) in the fluorescence detection system.
+A custom built lens assembly allows us to fix the first collection lens at about 30 mm from the crystal growth spot, enabling efficient collection of the emitted signal which is then focused on either a spectrometer or a photodiode.
 In this paper, we extend the approach conducted in previously cited works, proving the first successful growth of good quality Ba doped neon crystals at temperatures ranging from 6.8 to 9 K.
-Among the hosts previously studied, neon is a promising matrix for MIT, as it can offer good optical transparency and reduced perturbation of embedded species, as proven in previous studies of atoms embedded in cryogenic neon crystals 1Yb_Neon_Lambo, dargyte2021optical_Rb_Neon_Spin, braggio2022spectroscopy_RbNEon.
+Among the hosts previously studied, neon is a promising matrix for MIT, as it can offer good optical transparency and reduced perturbation of embedded species, as proven in previous studies of atoms embedded in cryogenic neon crystals Lambo et al.
+[2012], Dargyte et al.
+[2021], Braggio et al.
+[2022].
 Therefore, this approach enables direct comparison of Ba properties in our neon crystal lattice with respect to the previously studied ones.
-In this work, we investigate neutral barium atoms embedded in a neon matrix, where neon serves both as a buffer gas in our atomic source and as a matrix deposition gas for crystal growth.
-This systems enables LIF spectroscopic measurements of barium atoms in the crystal, using single and double laser excitation in either continuous or pulsed mode.
+In this work, neon serves both as a buffer gas in our atomic source and as a matrix deposition gas for crystal growth.
+Our setup permits LIF spectroscopic measurements of barium atoms in the crystal, using single and double laser excitation in either continuous or pulsed mode.
 These methods enable selective excitation of barium energy levels.
 We also measure, for the first time, the lifetime of the metastable state 5d6s 3D1 of barium in a neon matrix.
-This approach creates not only a new base for spectroscopic analysis of barium in inert cryogenic matrices, but also an essential experimental framework for embedding molecules, such as barium monofluoride (BaF), in cryogenic crystals.
-BaF is a molecule with great characteristics for low energy, high precision measurements such as the electron electric dipole moment (eEDM) through the MIT.
-BaF production will include deposition of some neutral barium presence in the matrix, unless it is filtered out yau2024specularreflectionpolarmolecules, corriveau2024matrixisolatedbariummonofluoride at the cost of a loss in the number of molecules.
-Therefore, understanding the spectral signatures and dynamics of Ba will be essential, as it will provide a crucial spectroscopic baseline for future BaF matrix isolation measurements in a cryogenic inert environment such as neon Li_2023, corriveau2024matrixisolatedbariummonofluoride or parahydrogen Messineo_pH, Borghesani_2023, as recently proposed by our group within the DOCET experiment.
+The study of barium in cryogenic inert matrices serves as an initial spectroscopic approach towards molecules, such as BaF, with great characteristics for low energy, high precision measurements such as the electron electric dipole moment (eEDM) through the MIT.
+BaF production will include deposition of some neutral barium presence in the matrix, unless it is filtered out Yau et al.
+[2024], Corriveau et al.
+[2024] at the cost of a loss in the number of molecules.
+Therefore, understanding the spectral signatures and dynamics of Ba will be essential, as it will provide a crucial spectroscopic baseline for future BaF matrix isolation measurements in a cryogenic inert environment such as neon Li et al.
+[2023], Corriveau et al.
+[2024] or parahydrogen Antonini et al.
+[2025], Borghesani et al.
+[2023], as recently proposed by our group within the DOCET experiment.
 II Experimental setup
@@
 II.1 Barium doped neon crystals production
-[…]
+Neutral barium atoms are produced in a custom designed stainless steel cryogenic ablation cell, place in a vacuum chamber maintained below $1\times 10^{-7}$ mBar as shown in Figure 1 A), exploiting the buffer gas cooling technique Doyle et al.
+[1995], Bethlem et al.
+[2002].
+Laser ablation, within this method, allows for the production of large numbers of atoms at very low temperature, as collisions with the cold buffer gas cool them down.
+In our setup, neon is used as a buffer gas.
+The cell is cooled with a pulse tube refrigerator Leybold-Heraeus RG210 and kept at approximately 20 K, to avoid neon condensation while still allowing efficient cooling of the produced atoms.
 Ablation laser pulses at 1064 nm with a repetition rate of 10 Hz and energies between 10 and 30 mJ, generated from a EKSPLA NL300 Nd:YAG laser, hit the metallic barium target, fixed on one side of the cell, on a focused spot with diameter of approximately 0.5 mm.
@@
 The ablation spot is periodically shifted to maintain stable production.
-To probe the production of neutral barium atoms in the gas phase, before deposition, two absorption measurements are performed simultaneously, as already done in previous studies for BaF mooij2024novel, inside and outside of the cell.
+To probe the production of neutral barium atoms in the gas phase, before deposition, two absorption measurements are performed simultaneously, as already done in previous studies for BaF Mooij et al.
+[2024], inside and outside of the cell.
 An Avesta Ti:Sa laser is used, with a linewidth of approximately 2 GHz, tunable over a 14 300 - 11100 cm-1 range (700 - 900 nm).
@@
 At a repetition of 10 Hz, the pulses are 10 ns long and on a range between 1 to 2.5 mJ and are used to excite neutral barium to high energy states, as describe in the next section.
-Light collection occurs almost orthogonally from the excitation beam, to reduce the unwanted excitation signal and optimize the fluorescence detection.
+Light collection occurs almost orthogonally from the excitation beam, as illustrated in Figure 1 in purple arrows, to reduce the unwanted excitation signal and optimize the fluorescence detection.
 The light detection system consisted of a two lens assembly to collect fluorescence emitted from the substrate.
@@
 The work of Lebedev et al.
-Weis_BaHelium demonstrate that a third harmonic of an Nd:YAG laser, with pulses at 28 195 cm-1 (355 nm) can excite a wide range of barium states by both radiative or non-radiative transitions in a solid helium matrix.
+Lebedev et al.
+[2011] demonstrate that a third harmonic of an Nd:YAG laser, with pulses at 28 195 cm-1 (355 nm) can excite a wide range of barium states by both radiative or non-radiative transitions in a solid helium matrix.
 Among these, long lived metastable state of barium, such as 5d6s 1D2 and the 5d6s 3Dj ($j=1,2,3$) manifold, act as an intermediate reservoir, and when a second pulse excites them, they can reach excited state at energies over 30 000 cm-1 with respect to the energy of the ground state.
 This method leads to barium fluorescence emission through a radiative cascade, as from the highest excited state the atom can populate a metastable state before decaying into the ground state.
-We adapted this approach to detect barium fluorescence in a neon crystal (vis-NIR), enabling the first comparison up to date with the work of Lebedev et al. Weis_BaHelium in solid helium.
+We adapted this approach to detect barium fluorescence in a neon crystal (vis-NIR), enabling the first comparison up to date with the work of Lebedev et al. Lebedev et al. [2011] in solid helium.
 As illustrated in Figure 3 and summarized in Table 1 we observe a broad set of barium fluorescence peaks when a third harmonic of an Nd:YAG pulsed laser is shined onto the crystal.
@@
 The spectral shifts are referenced against the corresponding free atomic barium lines, listed in the lower portion of the table.
-Transition H is an electric dipole forbidden line in the gas phase NIST_ASD.
+Transition H is an electric dipole forbidden line in the gas phase Kramida et al.
+[2024].
 Symbol $\ast$ = multiple transitions involved.
@@
 Comparing our data with the spectroscopic results of Lebedev et al.
-on barium in solid helium Weis_BaHelium, four fluorescence lines match those previously reported.
+on barium in solid helium Lebedev et al.
+[2011], four fluorescence lines match those previously reported.
 In general, the relatively narrow wavelength shifts with respect to the free barium atomic lines and the relatively small full width at half maximum (FWHM) allow us to reliably identify five electronic state transitions (C, E, F, G, H, see Figure 3), as shown in Table 1.
 Among these, the largest observed shift is for peak C (93-93 cm-1), while the smallest is for peak F (9-9 cm-1).
-The magnitude of these shifts is smaller than those reported for barium isolated in rare gas matrices McCaffrey.
+The magnitude of these shifts is smaller than those reported for barium isolated in rare gas matrices Davis and McCaffrey [2018].
 However, some emission features may not be fully resolved because of their relatively large FWHM.
@@
 The FWHM values can be directly compared with those obtained by Lebedev et al.
-Weis_BaHelium of barium in solid helium at 1.5 K.
+Lebedev et al.
+[2011] of barium in solid helium at 1.5 K.
 Peak C, corresponding to the 6s6p $^{1}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$ transition, has a width of $\Delta\nu=156\ \text{cm}^{-1}$, broader than the helium value ($\delta\nu=30\ \text{cm}^{-1}$).
@@
 In this section, our objective is to investigate the doped crystal using continuous-wave lasers, following the work of Li et al.
-Li_2023.
+Li et al.
+[2023].
 To this end, we irradiated 1 mm-wide spots on the crystal surface with Ti:Sa lasers delivering up to 100 mW, and measured excitation spectra by tuning the laser over its operational range of 14 300–11 100 cm-1 (700–870 nm).
@@
 Those states can decay to the ground state through radiative cascade involving the 5d6s 3Di (i = 1,2,3) and 1D2 states.
-Direct fluorescence from the 3Pj states is not detected, as it spectrally overlaps with the excitation light.
-Our working hypothesis is that the excitation laser directly addresses the 6s6p 3P2 level at 13 514.7 cm-1 (740.0 nm), a strongly suppressed transition in the free atom.
-However, in the matrix environment, the local perturbation breaks the atomic symmetries that govern the selection rules in the gas phase moroshkin2008atomic.
-As a result, transitions that are forbidden in free barium become weakly allowed, allowing the population transfer to the 6s6p 3P2 state and the subsequent relaxation into the 5d6s 3Di manifold or 5d6s 1D2 state.
-As stated in precedent spectroscopy studies of Ba atoms in RG gas matrices, the line broadening of excitation is larger with respect to the fluorescence emission.
+Direct fluorescence from the 3Pj states is not detected, as it might be too close to the excitation light which was properly filtered out.
+Stoke shifts are usually big enough to directly detect emission light, but in neon such transitions have not been detected in the spectral region of collection.
+In the matrix environment, the local perturbation breaks the atomic symmetries that govern the selection rules in the gas phase Moroshkin et al.
+[2008].
+As a result, transitions that are forbidden for free barium become weakly allowed.
+In our case, the excitation laser seems to directly address the 6s6p 3P2 level at 13 514.7 cm-1 (740.0 nm), a strongly suppressed transition in the free atom, allowing the population transfer to the 6s6p 3P2 state and the subsequent relaxation into the 5d6s 3Di manifold or 5d6s 1D2 state.
+As stated in precedent spectroscopy studies of Ba atoms in RG gas matrices Davis and McCaffrey [2018], the line broadening of excitation is larger with respect to the fluorescence emission.
 In this case, the effect is likely enhanced by the indirect excitation discussed above.
@@
 An emission band is observed around 9033.4 cm-1 (1107.1 nm), corresponding to population of the 5d6s 3D1 state, with maximum intensity under excitation at 13 514 cm-1 (749.8 nm) and long blue tail of approximately 40 nm, similar to the one observed previously in Figure 5.
-Interestingly, we also detected green fluorescence around 18000 cm-1 corresponding to the 6s6p 1P1 state, as shown in Figure 7 (a).
+Green fluorescence around 18000 cm-1 corresponding to the 6s6p 1P1 state has been detected, as shown in Figure 7 (a).
 The photon energies provided by our excitation laser are insufficient to directly excite the 6s6p 1P1 state.
@@
 The resulting spectrum, shown in Figure 7 (b), reveals an increase in the fluorescence intensity of the previously detected signal, as well as the appearance of new excitation combinations, resulting in high intensity fluorescence of the 6s6p 1P1 state.
-Interestingly the 6s6p 3P1 state exhibits two dominant fluorescence bands centered at 18 622 (557 nm) and 17 967 (538 nm) cm-1.
-Similar features have been extensively studied in Ar, Kr, and Xe matrices McCaffrey, where they are attributed to the influence of lattice site occupancy on barium emission properties.
+The 6s6p 3P1 state exhibits two dominant fluorescence bands centered at 18 622 (557 nm) and 17 967 (538 nm) cm-1.
+Similar features have been extensively studied in Ar, Kr, and Xe matrices Davis and McCaffrey [2018], where they are attributed to the influence of lattice site occupancy on barium emission properties.
+Further investigation of the process of site selection of Barium in a neon crystal is out of the scope of our project and has already been discussed for BaF embedded in a neon matrix Lambo et al.
+[2023].
 Figure 8: Energy level diagram (not to scale) of the proposed double laser pumping scheme for barium in a neon matrix.
@@
 From this level, a second laser (blue arrow) tuned at 12 849 cm-1 (780 nm) excites the atom to the higher 5d6p 3F2 state.
-According to the cascade dynamics reported in previous works Weis_BaHelium, this state decays to the 5d6s 1D2 level with emission at 10 670 cm-1 (937 nm), followed by a second fluorescence decay to the ground state 6s2 1S0 at 11 394 cm-1 (877 nm).
+According to the cascade dynamics reported in previous works Lebedev et al.
+[2011], this state decays to the 5d6s 1D2 level with emission at 10 670 cm-1 (937 nm), followed by a second fluorescence decay to the ground state 6s2 1S0 at 11 394 cm-1 (877 nm).
 The black arrows represent spontaneous emission pathways, while red arrows denote laser driven excitations.
 Dashed lines indicate indirect population routes from the 6s6p 3Pj manifold.
-We observe a high intensity fluorescence emission when the excitation laser is tuned to approximately 14 180 cm-1 (705 nm).
-However, it does not correspond to any neutral atomic barium state, and no direct excitation pathway from the ground or low lying metastable states to higher lying electronic states has been identified at this wavelength.
-This discrepancy suggests that the observed fluorescence may arise from matrix induced perturbations, site selective effects, or indirect population mechanisms through long lived intermediate states.
-Further investigation is beyond the scope of this work.
+We observe a high intensity fluorescence emission when the excitation laser is tuned to approximately 14 180 cm-1 (705 nm) which is probably populating the 5d6s 1D2 state as shown in Figure 5, as well as two dominant emission when a second laser excites the sample at around 12 270 cm-1 (815 nm) and 12 060 cm-1 (830 nm).
+We note that the three excitations are close to BaF energy levels reported in previous works Li et al.
+[2023].
+The observed emission is originating from different population pathways involving long living intermediate states, as shown in the following sections.
 To further test the feasibility of the double–laser excitation, we designed a pumping scheme path that populates 5d6p 3F2 at 22 065 cm-1 (453 nm) only if we strongly populate the manifold 6s6d $^{3}\text{D}_{j}$, as shown in Figure 8.
 Although the excitation bands are $\approx$1000 cm-1 wide, our measurements demonstrate that this scheme works exclusively via indirect population of the first manifold, specifically through the 5d6s 3D2 state.
-A second laser, set at 12 849 cm-1, can then couple this state and populate the 5d6p 3F2 state laying at 22 064.622\,064.6 cm-1 (453 nm).
+A second laser, set at 12 849 cm-1, can then couple this state and populate the 5d6p 3F2 state laying at 22 06522\,065 cm-1 (453 nm).
 This scheme relays on the evidence shown earlier that excitation near 13 369 cm-1 (748 nm) populates the 5d6s 3D1 and hypothetically the other states of the same 5d6s 3Dj manifold.
@@
 To further test the scheme, if we tune the first laser to 13 504 cm-1 (740 nm), a known population maximum of the 5d6s 1D2 state, no excitation pathway to the 5d6p 3F2 level is observed, as shown in Figure 9 (b), probing our selective pumping scheme functioning.
+Moreover, both lasers have been applied to the crystal at higher intensities, individually and simultaneously, as saturation is usually never reached in a matrix isolation environment and additional atom emissions might arise.
+However, the obtained spectra exhibit the same characteristic as those reported here.
 III.3 Lifetime of first excited state 5d6s 3D1
 As shown in the previous sections, and schematically in Figure 4, a single laser allows us to populate the 5d6s 3Dj manifold, similar to the earlier work of Lebedev et al.
-Weis_BaHelium, where the 5d6s 3D1 state of barium in a helium matrix at 1.5 K is populated using the pulsed second harmonic of a Nd:YAG laser at approximately 18 80018\,800 cm-1 .
-In our experiments with barium in a solid neon at 6.8 K, this state is populated using continuous wave laser light tuned to 13 514 cm-1, enabling a direct comparison with the lifetime measurement performed in solid helium and in the vacuum gas phase.
+Lebedev et al.
+[2011], where the 5d6s 3D1 state of barium in a helium matrix at 1.5 K is populated using the pulsed second harmonic of a Nd:YAG laser at approximately 18 80018\,800 cm-1 .
+In our experiments with barium in a solid neon at 6.8 K, this state is populated using continuous wave laser light tuned to 13 514 cm-1, enabling a direct comparison with the lifetime measurement performed in solid helium.
 The lifetime measurement is carried out using an Electro Optic Modulator (EOM), driven by a pulse generator.
@@
 At temperature higher than 8 K, intensity in the fluorescence shows a plateau.
-We obtain a lifetime more than two orders of magnitude higher than the gas phase measurement, which is estimated to be 1.5 $\pm$ 0.2 ms AdamsLeiHusain+2001+1205_lifetimeBaGas.
-This increase might indicate the weak interaction of the host matrix, as neon low phonon density and weak interaction with the embedded atoms enable long lived excited states.
-For a comparison, in solid helium, an even better lifetime of $2.72\pm 0.04$ s is observed at a temperature of 1.5 K Weis_BaHelium.
+For a comparison, in solid helium, a longer lifetime of $2.72\pm 0.04$ s is observed at a temperature of 1.5 K Lebedev et al. [2011].
 Lifetime measurements have been conducted at different temperatures, as, ranging from 6.8 to 8 K, due to temperature limitations of the setup.
@@
 A rate equation model has been used to fit the obtained lifetime and extrapolate the possible lifetime values at lower temperatures.
-It considers both radiative ($\tau_{r}$) and non-radiative ($\tau_{nr}$) decay channels, described by the two termsfit_description_PL, torVergata_fitArrhenius:
+It considers both radiative ($\tau_{r}$) and non-radiative ($\tau_{nr}$) decay channels, described by the two termsKobitski et al.
+[2003], Prestopino et al.
+[2019]:
 $\frac{1}{\tau(T)}\;=\;\frac{1}{\tau_{r}}\;+\;\frac{1}{\tau_{nr}}\,\exp\!\left(-\frac{E_{a}}{k_{B}T}\right)\,.$ (1)
-Fit results are shown in Table 2, where $E_{a}$ is the activation energy . The lifetime is expected to increase to approximately 0.42 $\pm$ 0.03 s at 2 K.
+The fit results are shown in Table 2, where $E_{a}$ is the activation energy.
+The lifetime is expected to increase to approximately 0.42 $\pm$ 0.03 s at 2 K.
+T The model captures the expected low temperature plateau corresponding to a regime where the dopant non radiative decay channels are suppressed.
+Moreover, in the precedent work of BaF in neonLi et al.
+[2023], a similar model has been used, with a resulting activation energy value of $E_{a(BaF:Ne)}=hc\times 16.2(3)\;\text{cm${}^{-1}$}=(2.01\pm 0.04)\times 10^{-3}\;\text{eV}$, consistent with the one obtained in this work, shown in shown in Table 2.
+Future work and new models might take into account additional mechanisms inherent to matrix environments, such as site dependent trapping, crystal inhomogeneity and local stress of laser excitation light.
 Table 2: Fit parameters from the rate equation model describing the temperature dependent lifetime of the 5d6s 3D1 state of barium in solid neon.
@@
 IV Conclusions
-For the time in literature, we observed barium atoms embedded in a neon matrix through a combination of cascade florescence and selective excitation using a third harmonic 1064 nm pulsed laser and single or double continuous wave lasers, exploiting long living metastable state of barium within the matrix.
+For the first time in literature, we observed barium atoms embedded in a neon matrix through a combination of cascade fluorescence and selective excitation using a third harmonic 1064 nm pulsed laser and single or double continuous wave lasers, exploiting long living metastable state of barium within the matrix.
 The observed fluorescence spectra display small shifts relative to the gas phase transitions and moderate inhomogeneous broadening, indicating the weak perturbation exerted by the neon matrix.
@@
 Previous work by Lancaster et al.
-Lancaster_2021_NevsPh found no detectable fluorescence for rubidium in parahydrogen, whereas clear emission was observed for boron Tam2000_boron_pH and other molecules Huang2016_molecules_pH.
+Lancaster et al.
+[2021] found no detectable fluorescence for rubidium in parahydrogen, whereas clear emission was observed for boron Tam et al.
+[2000] and other molecules Huang and Lee [2016].
 A similar investigation for Ba and BaF will be crucial to establish parahydrogen as a next-generation host for matrix-isolated spectroscopy and precision measurements of fundamental properties.